OpenStudy (anonymous):

solve please 1/4(x/3-2) =-3/4

OpenStudy (anonymous):

Where is the(x/3 -2) - above or below the fraction line?

OpenStudy (anonymous):

x over 3 -2

OpenStudy (anonymous):

\[\frac{ 1 }{ 4(x/3 - 2)} \]

OpenStudy (anonymous):

Like this ?

OpenStudy (anonymous):

Or like this\[\frac{ 1 }{ 4 }(x/3 -2)\]

OpenStudy (anonymous):

nope its just 1/4 (x/3-2/1) = -3/4

OpenStudy (anonymous):

Omg - please attach a clipping of the exact expression (use "print screen") or smthng

OpenStudy (anonymous):



OpenStudy (anonymous):

U have a common denominator 4 ==> multiply both sides by it